can you help me solve 5m-n=11 and3m+2n=4
Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

5m-n=11 --- Eqn 1
3m+2n=4 --- Eqn 2
==================
Multiply Eqn 1 by 2: ---> 10m-2n = 22
---
3m + 2n = 4 --- Eqn 2
10m -2n = 22 -- Eqn 1 times 2
---------------------------------------- Add
13m = 26
m = 2
etc

Answer by ikleyn(52781)

.

It looks like you need to learn  (to get familiar with)  the general theory/technique on solving systems of two equations in two unknown.

In this site I developed such lessons for beginner students.  These lessons are
    - Solution of the linear system of two equations in two unknowns by the Substitution method
    - Solution of the linear system of two equations in two unknowns by the Elimination method
    - Solution of the linear system of two equations in two unknowns using determinant
    - Geometric interpretation of the linear system of two equations in two unknowns
    - Useful tricks when solving systems of 2 equations in 2 unknowns by the Substitution method

They give you the general theory and a lot of examples which you can consider as your TEMPLATES.

Consider these lessons as your textbook,  your handbook,  your guide and your  (free of charge)  home teacher.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Systems of two linear equations in two unknowns".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.